### v1

# Singular Value Scaling: Efficient Generative Model Compression via Pruned Weights Refinement

Abstract
While pruning methods effectively maintain model performance without extra training costs, they often focus solely on preserving crucial connections, overlooking the impact of pruned weights on subsequent fine-tuning or distillation, leading to inefficiencies. Moreover, most compression techniques for generative models have been developed primarily for GANs, tailored to specific architectures like StyleGAN, and research into compressing Diffusion models has just begun. Even more, these methods are often applicable only to GANs or Diffusion models, highlighting the need for approaches that work across both model types. In this paper, we introduce Singular Value Scaling (SVS), a versatile technique for refining pruned weights, applicable to both model types. Our analysis reveals that pruned weights often exhibit dominant singular vectors, hindering fine-tuning efficiency and leading to suboptimal performance compared to random initialization. Our method enhances weight initialization by minimizing the disparities between singular values of pruned weights, thereby improving the fine-tuning process. This approach not only guides the compressed model toward superior solutions but also significantly speeds up fine-tuning. Extensive experiments on StyleGAN2, StyleGAN3 and DDPM demonstrate that SVS improves compression performance across model types without additional training costs. Our code is available at: https://github.com/LAIT-CVLab/Singular˙Value˙Scaling.
Introduction
Generative models like Generative Adversarial Networks (GANs) (Goodfellow et al. 2014) and Diffusion models (Ho, Jain, and Abbeel 2020) have achieved remarkable performance across various computer vision tasks such as image generation (Karras et al. 2020; Ho, Jain, and Abbeel 2020; Wang et al. 2022), image editing (Baykal et al. 2023; Pehlivan, Dalva, and Dundar 2023; Kawar et al. 2023; Zhang et al. 2023), even video generation (Skorokhodov, Tulyakov, and Elhoseiny 2022; Ho et al. 2022; Blattmann et al. 2023) and 3D generation (Chan et al. 2022; Karnewar et al. 2023). The impressive performance of these generative models, however, often comes at the cost of high memory and computational demands, limiting their real-world applicability. To address these issues, several model compression technique for generative models have been proposed (Liu et al. 2021; Xu et al. 2022; Chung et al. 2024; Fang, Ma, and Wang 2023).
Model compression typically involves two steps: 1) pruning to reduce model size while retaining essential pre-trained knowledge, and 2) fine-tuning to restore performance. Among these, effective pruning has received significant attention for its ability to enhance following fine-tuning step. The retained pre-trained knowledge from pruning can enhance fine-tuning, leading to improved performance and faster convergence without additional training costs. While previous pruning methods (Chung et al. 2024; Fang, Ma, and Wang 2023) effectively maintain model performance, they often focus solely on preserving crucial connections in the pre-trained model, overlooking the impact of pruned weights on subsequent process, leading to inefficient fine-tuning or distillation. This issue becomes more severe as the model’s capacity decreases, and in some cases, the pruned weights results in worse performance compared to random initialization; i.e. slow convergence speed and lower performance. Therefore, addressing these factors is essential for achieving more efficient generative models.
Our key observation is that pruned weights often exhibit dominant singular vectors, which results in a large disparity between the largest and smallest singular values. The presence of such dominant singular vectors significantly impacts the model’s forward propagation, overshadowing the contributions of minor singular vectors, and the overall fine-tuning process is dominated by these few dominant singular vectors. We find that this could limit the exploration of diverse weight space. Therefore, ensuring a balanced contribution among the singular vectors within pruned weights at initialization may offer a potential solution to the inefficiencies observed in the fine-tuning process.
In this paper, we introduce a simple yet effective refinement technique called Singular Value Scaling (SVS) to enhance the efficiency of fine-tuning pruned weights. The dominant singular values is scaled down to reduce the disparity between the largest singular values and smallest singular values while preserving the relative order of the singular values at initialization. In this way, the refined pruned weights makes the fine-tuning process easier compared to directly using the pruned weights by ensuring all singular vectors contribute more evenly at the beginning of fine-tuning. Note that since our method focuses on the knowledge within each weight and is independent of specific architectures, it can be applied to both GANs and diffusion models.
We conduct extensive experiments with representative generative model architectures across various datasets, including StyleGAN2, StyleGAN3, and Denoising Diffusion Probabilistic Model (DDPM) on CIFAR10, CelebA-HQ, FFHQ, and LSUN-Church. The results demonstrate that our method enhances the fine-tuning process, leading to both faster convergence and improved solutions for the compressed model without additional training cost. Our contributions can be summarized as follows:
  • We propose a simple yet effective method that enhances the efficacy of fine-tuning process of pruned generative models without additional training cost.
  • By simply scaling down the singular values of pruned weights at initialization, our method, Singular Value Scaling (SVS) enables the compressed model to converge faster and achieve superior performance than using the existing baseline methods.
  • We are the first to provide a general method for generative model compression that can be applied to both model types of GANs and Diffusion models.
Related Works
StyleGAN Compression
Recently, several studies (Liu et al. 2021; Xu et al. 2022; Chung et al. 2024) have been proposed for compressing unconditional GANs, particularly the StyleGAN family (Karras et al. 2020, 2021), which are the state-of-the-art models in the area. CAGAN (Liu et al. 2021) introduced the first framework for pruning pre-trained StyleGAN models and fine-tuning them via pixel-level and feature-level knowledge distillation. Following this, StyleKD (Xu et al. 2022) further enhanced the fine-tuning process by incorporating a specialized relation loss tailored for StyleGAN. In particular, CAGAN introduced a content-aware pruning technique that preserves connections in the pre-trained model crucial for the semantic part of generated images, while StyleKD tackled output discrepancies problem by inheriting only the pre-trained model’s mapping network and randomly initializing the synthesis network with smaller size. More recently, DCP-GAN (Chung et al. 2024) proposed a diversity-aware channel pruning technique, which preserves connections in the synthesis network that contribute to sample diversity, significantly enhancing both the diversity of generated images and the training speed.
Diffusion Model Compression
Diffusion models have garnered significant attention for their stable training and impressive generative capabilities. However, their performance incurs high computational costs due to iterative sampling. To enhance efficiency, various sampling techniques have been developed to reduce the number of required iterations (Song, Meng, and Ermon 2021; Lu et al. 2022; Zheng et al. 2023). Orthogonal to the these sampling techniques, Diff-Prune (Fang, Ma, and Wang 2023) introduced a seminal work for reducing computational costs by compressing Diffusion models. This demonstrated that by primarily pruning connections involved in less important diffusion steps, a reduced-size Diffusion model with minimal performance loss can be achieved in significantly fewer training iterations compared to the original-sized model.
Weight Initialization and Trainability
Weight initialization is essential for efficient training in deep learning. Glorot (Glorot and Bengio 2010) and He initialization (He et al. 2015) are widely used to maintain activation variance and ensure stable optimization, particularly in feed-forward networks with ReLU activations. Orthogonal initialization (Saxe, McClelland, and Ganguli 2014), which ensures that all singular values of weights are 1, achieves dynamical isometry, the ideal state for trainability. In classifier pruning, TPP (Wang and Fu 2023) recently highlighted the importance of preserving trainability in pruned networks to support effective performance recovery.
Analysis
Figure 1: (a) A schematic overview of our compression framework. Unlike existing model compression scheme, we additionally perform pruned weights refining step for efficient model fine-tuning. First, we prune the pre-trained model. Next, we compute singular values of the weights of the pruned model. Then, we refine the pruned weights by scaling down the singular values with large magnitudes and scaling up the singular values with small magnitudes (Singular Value Scaling, SVS). Finally, the refined pruned model are fine-tuned. (b), (c) The singular value distribution before and after singular value scaling in the pruned weights by the method of DCP-GAN and Diff-Prune. The x-axis represents singular values and the y-axis represents the density. When applied SVS, the singular value disparity reduces, balancing the power of each singular vector in the pruned weights.
Figure 2: FID convergence graph in different StyleGAN2 Architectures compressed by different methods. The x-axis represents the number of images shown to the discriminator. Solid line represents FID with respect to the training images, while dashed line represents the best FID of the compressed model by the corresponding method. “scaled” means that our Singular Value Scaling is applied to the pruned weights.
suboptimal Results with Pruned Weights
Previous works in the field of generative model compression (Chung et al. 2024; Fang, Ma, and Wang 2023) have shown that well-designed pruning techniques can lead compressed models to better solutions compared to their randomly initialized counterparts, while reaching the same performance more quickly (×$\times 2.5$ for StyleGAN2, ×$\times 8.0$ for DDPM). This is because pruned weights provide the compressed model with an initialization state that preserves the pre-trained model’s performance, starting with better performance compared to the random counterpart. This initial gain continues throughout the fine-tuning process, ultimately resulting in a better-performing compressed model. However, we observe that as fine-tuning progresses, models initialized with pruned weights converge more slowly compared to those initialized randomly (see Figure 2: orange line (DCP-GAN) vs. blue line (StyleKD)). This issue worsens as the model size decreases, resulting in suboptimal solutions, which indicates that the pruned weights themselves contain factors that hinder fine-tuning efficiency.
Analyzing Pruned Weights with SVD
We analyze the learned prior of pruned weights inherited from the pre-trained model by employing Singular Value Decomposition (SVD), which is widely adopted for low-rank approximation by identifying important basis from the weights (Denton et al. 2014; Zhang et al. 2015; Girshick 2015; Yoo et al. 2019). For a weight matrix $W\in\mathcal{R}^{m\times n}\ (m<n)$, we can decompose it using SVD:
$W=U\Sigma V^{T}=\sum^{m}_{i=1}\sigma_{i}\vec{u}_{i}\vec{v}_{i}^{T}$
[…]
Method
Scaling Singular Values of Pruned Weights
Based on our observation, we propose “Singular Value Scaling (SVS)”, to refine pruned weights to enhance fine-tuning efficiency. Our primary goal is to reduce the gap between the singular values of pruned weights. In the pruned weights, dominant singular vectors tend to have significantly larger singular values compared to smaller ones, and this gap increases as the values grow. Since all bases in pruned weights contain important knowledge from the pre-trained model, we prevent any single basis from dominating to let these bases contribute equally at the beginning of training. To achieve this, we simply scale the singular values using the “square root function”,
$W_{scaled}=U\Sigma_{scaled}V^{T}=\sum^{m}_{i=1}\sqrt{\sigma_{i}}\vec{u}_{i}\vec{v}_{i}^{T}$
Here, $U$ and $V$ remain unchanged, and $\Sigma_{scaled}=diag(\sqrt{\sigma_{1}},\sqrt{\sigma_{2}},\cdots,\sqrt{\sigma_{m}})\in\mathcal{R}^{m\times n}$. Square root function has several good properties: 1) Since singular values are non-negative, it maps non-negative values to non-negative values, 2) For $\sigma_{i}<1$, it increases the value, while for $\sigma_{i}>1$, it decreases the value more as the value grows. 3) In the positive domain, it is monotonically increasing, thereby maintaining the relative order of singular values. By scaling the singular values of pruned weights in this manner, we can preserve the original bases while balancing their relative contributions. This balanced contribution helps the compressed model fully leverage the pre-trained model’s knowledge, facilitating a more effective path to the optimal solution. Alternative functions, such as $\log(x+1)$ or $|\log(x)|$, could also achieve similar effects. We provide ablation study on these functions in the Experiments section.
Scaling Bias with respect to the Scaled Weights
Since biases and weights are learned simultaneously, we need to consider the impact of scaling the singular values of the weights on the corresponding biases. Let us consider the following linear equation, $y=Wx+b,$ where $W\in\mathcal{R}^{m\times n}$ is the weights, $x\in\mathcal{R}^{n}$ is the input and $b\in\mathcal{R}^{m}$ is the bias. We can factor the equation with respect to $W$ as follows:
$y=W(x+W^{\dagger}b)=W(x+V\Sigma^{-1}U^{T}b)$
where $W^{\dagger}\in\mathcal{R}^{n\times m}$ is the Moore-Penrose inverse (Petersen 2012) and $\Sigma^{-1}=diag(\sigma^{+}_{1},\sigma^{+}_{2},\cdots,\sigma^{+}_{m})\in\mathcal{R}^{n\times m}$ with
$\sigma^{+}_{i}=\begin{cases}\frac{1}{\sigma_{i}},&\text{if }\sigma_{i}\neq 0\\ 0,&\text{if }\sigma_{i}=0\end{cases}$
By representing the bias $b$ as $b=|b|\vec{\beta}$, where ||$\cdot$|| ⋅ | is the vector norm and $\vec{\beta}$ is the normalized vector of $b$, it becomes:
$y=W(x+V\Sigma^{-1}U^{T}b)=W(x+V(|b|\Sigma^{-1})U^{T}\vec{\beta})$
where $|b|\Sigma^{-1}=diag(|b|\sigma^{+}_{1},|b|\sigma^{+}_{2},\cdots,|b|\sigma^{+}_{m})$. Thus, for $\sigma_{i}>0$, when scaling the biases, instead of only scaling $\sigma_{i}$, we must also scale $b$: $b_{scaled}=b/\sqrt{|b|},$ which leads to:
$|b_{scaled}|/\sigma_{i,scaled}=\sqrt{|b|/\sigma_{i}}.$
Experiments
Dataset Method Arch Params $\downarrow$ FLOPs $\downarrow$ FID $\downarrow$ P $\uparrow$ R $\uparrow$ D $\uparrow$ C $\uparrow$
FFHQ Teacher base 30.0M 45.1B 4.29 0.762 0.561 1.061 0.845
StyleKD (Xu et al. 2022) 5.6M 4.1B 6.70 0.711 0.551 0.879 0.786
DCP-GAN (Chung et al. 2024) 6.51 0.712 0.556 0.884 0.785
DCP-GAN (scaled) 5.68 0.740 0.530 0.981 0.806
Teacher small 24.7M 14.9B 4.02 0.769 0.555 1.095 0.854
GS (Wang et al. 2020) 4.9M 1.3B 10.23 0.702 0.430 0.845 0.721
CAGAN (Liu et al. 2021) 9.23 0.685 0.500 0.760 0.722
StyleKD (Xu et al. 2022) 7.01 0.707 0.543 0.900 0.783
DCP-GAN (Chung et al. 2024) 8.17 0.680 0.539 0.787 0.741
DCP-GAN (scaled) 6.60 0.719 0.532 0.874 0.778
Church Teacher base 30.0M 45.1B 3.97 0.698 0.553 0.849 0.823
StyleKD (Xu et al. 2022) 5.6M 4.1B 5.72 0.703 0.460 0.845 0.777
DCP-GAN (Chung et al. 2024) 4.80 0.692 0.505 0.810 0.790
DCP-GAN (scaled) 4.62 0.716 0.501 0.878 0.812
Teacher small 24.7M 14.9B 4.37 0.696 0.512 0.815 0.809
StyleKD (Xu et al. 2022) 1.38B 1.38B 5.99 0.698 0.446 0.848 0.777
DCP-GAN (Chung et al. 2024) 6.30 0.687 0.435 0.793 0.762
DCP-GAN (scaled) 5.18 0.719 0.464 0.934 0.805
Table 1: Quantitative Results on StyleGAN2 compression. We report training results of previous methods on FFHQ-256 and LSUN Church-256 datasets. “Params” and “FLOPs” refer to the number of parameters in generator and floating-point operations respectively. “P” and “R” denote precision and recall metrics, while “D” and “C” represent density and coverage metrics. “base” and “small” denote the StyleGAN2 architecture with different capacity. “scaled” means that our Singular Value Scaling is applied to the pruned weights. All reported metrics are trained results with the official StyleGAN2 implementation.
Method FID $\downarrow$ P $\uparrow$ R $\uparrow$ D $\uparrow$ C $\uparrow$
Teacher 4.76 0.736 0.593 0.955 0.817
DCP-GAN 9.88 0.694 0.484 0.780 0.720
DCP-GAN (scaled) 8.38 0.702 0.531 0.808 0.748
Table 2: Quantitative Results on StyleGAN3 Compression.
DDPM CIFAR-10 32$\times$32 (100 DDIM steps)
Method Channel Sparsity $\uparrow$ Params $\downarrow$ MACs $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ P $\uparrow$ R $\uparrow$ D $\uparrow$ C $\uparrow$ Train Steps $\downarrow$
pre-trained 0% 35.7M 6.1G 4.19 1.000 0.672 0.759 0.762 0.897 800K
Diff-Prune 30% 19.8M 3.4G 5.29 0.932 N/A N/A N/A N/A 100K
Diff-Prune 5.49 0.931 0.671 0.742 0.772 0.888 160K
Diff-Prune (scaled) 5.14 0.923 0.668 0.740 0.768 0.889
Diff-Prune 50% 8.96M 1.5G 7.77 0.914 0.670 0.720 0.805 0.861 360K
Diff-Prune (scaled) 7.41 0.912 0.671 0.718 0.817 0.868
Diff-Prune 70% 5.12M 1.0G 9.87 0.906 0.665 0.702 0.818 0.836 600K
Diff-Prune (scaled) 9.38 0.907 0.671 0.699 0.834 0.844
DDPM CelebA-HQ 64 × 64 (100 DDIM steps)
Method Channel Sparsity $\uparrow$ Params $\downarrow$ MACs $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ P $\uparrow$ R $\uparrow$ D $\uparrow$ C $\uparrow$ Train Steps $\downarrow$
pre-trained 0% 78.7M 23.9G 6.48 1.000 0.587 0.812 0.488 0.755 500K
Diff-Prune 30% 43.7M 13.3G 6.24 0.885 N/A N/A N/A N/A 100K
Diff-Prune 6.17 0.949 0.591 0.803 0.502 0.760 100K
Diff-Prune (scaled) 5.57 0.938 0.619 0.793 0.573 0.785
Diff-Prune 50% 19.7M 6.0G 5.32 0.926 0.618 0.773 0.594 0.793 200K
Diff-Prune (scaled) 4.66 0.921 0.634 0.766 0.626 0.812
Diff-Prune 70% 9.25M 4.2G 6.30 0.913 0.584 0.772 0.524 0.761 240K
Diff-Prune (scaled) 5.04 0.905 0.625 0.752 0.631 0.812
Table 3: Quantitative results on DDPM compression. $\dagger$” represents reported results in the paper. “N/A” indicates the inability to evaluate due to the inaccessibility of the weights used in the paper. “Channel Sparsity” refers the ratio of removed channels from the pre-trained model and “Params” and “MACs” are resulting number of parameters and multiply-accumulate of the compressed model. “SSIM” estimates the similarity between generated sample of pre-trained and compressed models given same noise. “P” and “R” denote precision and recall metrics, while “D” and “C” represent density and coverage metrics. “scaled” means that our Singular Value Scaling is applied to the pruned weights.
Experimental Setup
Baselines.
To evaluate our method, we conducted experiments on StyleGAN2, which is the most extensively studied model in the field of generative model compression. Additionally, to validate our method on generative models with different architectures, we conducted experiments on StyleGAN3 and the Denoising Diffusion Probabilistic Model (DDPM), which, along with StyleGAN2, are representative models in the field of generative modeling.
Implementation details.
For the StyleGAN2 compression, we re-implement previous StyleGAN2 compression methods, which were implemented on the unofficial StyleGAN2 implementation [^1] https://github.com/rosinality/stylegan2-pytorch, on the official StyleGAN2 implementation [^2] https://github.com/NVlabs/stylegan2-ada-pytorch because the official implementation provides more optimized StyleGAN2 training settings and are easy to use. we apply our method to the weights pruned by DCP-GAN (Chung et al. 2024), which is the state-of-the-art pruning method in StyleGAN2 compression. We mainly compare our method against StyleKD (Xu et al. 2022) and DCP-GAN because they are the methods that demonstrate the best performance in StyleGAN2 compression. We use two StyleGAN2 architectures: the optimized architecture provided by the official implementation (denoted as “StyleGAN2 (small)”) and the original structure used in previous works (denoted as “StyleGAN2 (base)”). Experiments are conducted on the FFHQ (Karras, Laine, and Aila 2019) and LSUN Church (Yu et al. 2015) datasets. Following DCP-GAN, we use a facial mask as the content mask for the FFHQ dataset and a uniform mask (all pixel values are assigned values 1) for LSUN Church dataset. For StyleGAN3 compression, we implement our method based on the DCP-GAN implementation [^3] https://github.com/jiwoogit/DCP-GAN/tree/stylegan3. Similar to StyleGAN2 compression, we apply our method to the pruned weights using the DCP-GAN method for our experiments. We follow the experimental setup of DCP-GAN and train StyleGAN3 on the FFHQ dataset until the discriminator see 10 million images for both DCP-GAN and our method. For DDPM compression, we build upon Diff-Prune (Fang, Ma, and Wang 2023) [^4] https://github.com/VainF/Diff-Pruning, a foundational work in Diffusion model pruning. We train pruned DDPMs using the CIFAR10 (Alex 2009) and CelebA-HQ (Liu et al. 2015) datasets, following the experimental setup of Diff-Prune. Unlike Diff-Prune, which tested only on a 30% channel sparsity, we explore more extreme levels of sparsity in our experiments, including 50% and 70% channel sparsity. We provide more detailed implementation in the supplementary material.
Evaluation metrics.
For StyleGAN compression, we evaluate our method using the Fréchet Inception Distance (FID) (Heusel et al. 2017) and Precision and Recall (P&R) (Kynkäänniemi et al. 2019). We also measure Density and Coverage (D&C) (Naeem et al. 2020), which are more robust to outliers than Precision and Recall. Here, Precision and Density evaluate the fidelity of generated samples, while Recall and Coverage assess their diversity. To calculate FID, we use all real samples from each dataset and 50K fake samples. For P&R and D&C calculation, we use 50K fake samples and all real samples from FFHQ, and 50K real samples from LSUN Church due to computational costs. For DDPM compression, in addition to FID, we employ Structural Similarity (SSIM) (Wang et al. 2004) to evaluate consistency with the pre-trained model, following the Diff-Prune. The SSIM score is measured between images generated by the pre-trained model and the compressed model, given the identical noise inputs. We use a 100-step DDIM sampler (Song, Meng, and Ermon 2021) for sampling.
Analysis on Convergence Speed
Figure 2 visualizes the FID convergence graph for two different StyleGAN2 architectures (StyleGAN2 (base) and StyleGAN2 (small)) in the FFHQ dataset. As shown in Figure 2 (a) (StyleGAN2 (base)), although DCP-GAN converges to better performance than StyleKD, the FID convergence speed of DCP-GAN noticeably slows down in the later stages of training. This issue becomes even more severe as the model’s capacity decreases. In Figure 2 (b), this slowdown allows StyleKD to surpass DCP-GAN early in fine-tuning, leading DCP-GAN to a suboptimal point. These results indicate that fine-tuning a model naively initialized with the pruned weights can lead to significantly poor outcomes, depending on the model’s capacity. In contrast, when pruned weights are refined using SVS, the model is much easier to fine-tune. It not only fits faster across most intervals but also converges to a significantly better solution compared to previous methods.
Quantitative Results on StyleGAN2 compression
Table 1 shows results on StyleGAN2 compression on different StyleGAN2 architectures in FFHQ and LSUN Church datasets. For FID, our method outperforms previous compression methods with a clear margin in all experiments. It generally outperforms previous methods in Precision and Density metrics, and matches or exceeds their performance in Recall and Coverage metrics. These results show that our method enables the compressed model to better match the distribution compared to previous methods, leading to higher quality and more diverse samples. One notable observation is that when the capacity of StyleGAN2 is sufficient (base), DCP-GAN consistently outperforms StyleKD, and as the capacity decreases (small), DCP-GAN’s performance lags behind that of StyleKD, indicating that the pruned weights from DCP-GAN are harder to fine-tune. In contrast, our method provides a more effective and refined initialization, resulting in improved final performance.
Quantitative Results on StyleGAN3 compression
Table 2 presents the results of StyleGAN3 compression. The model initialized with our refined pruned weights outperforms the DCP-GAN across all metrics. This result shows that, similar to StyleGAN2, refining pruned weights facilitates easier fine-tuning of the model, enabling it to generate more plausible and diverse images.
Quantitative Results on DDPM compression
Table 3 presents the results of DDPM compression. For DDPM compression, we compare our method with the best-performing baseline during fine-tuning. Diff-Prune achieves higher SSIM scores, indicating closer alignment with the pre-trained model. This is expected, as our method refines pruned weights to balance knowledge preservation with easier fine-tuning. Consequently, our method consistently achieves better FID scores across all pruning ratios, indicating more effective distribution matching. Additionally, the modest drop in SSIM indicates that our method still effectively retains contextual knowledge, even as it further relaxes the knowledge of the pre-trained model.
Figure 3: Qualitative results on FFHQ and LSUN Church datasets. Samples in each row are generated from same noise vector with StyleGAN2 (small), which is compressed using different compression methods with channel sparsity 70%. “Ours” denotes the compressed model with DCP-GAN refined by Singular Value Scaling.
Qualitative Results on StyleGAN2 compression
Figure 3 shows samples generated by StyleGAN2 (small) compressed using different methods from the same noise vector. Despite all compressed models being trained with the same loss, the characteristics of the generated samples vary depending on the weight initialization and frequently exhibit artifacts. For example, the model compressed by StyleKD often misses contextual details, such as face shape (second and fourth row). DCP-GAN, which prunes the pre-trained model to maximize diversity, tends to focus heavily on details like wrinkles and hair, while often failing to restore the overall appearance of the face. Additionally, as seen in the third column, first row of LSUN Church, DCP-GAN tends to generate excessively repetitive patterns by focusing too much on restoring fine details, often neglecting the overall context. In contrast, models with refined pruned weights preserve the details of the pre-trained model while faithfully restoring its overall context. This result shows that, to achieve both diversity and fidelity in the compressed model, it is essential not only to preserve knowledge from the pre-trained model but also to refine the pruned weights for a proper balance with fine-tuning.
Ablation Study on Other Pruning Method
Method FID $\downarrow$ P $\uparrow$ R $\uparrow$ D $\uparrow$ C $\uparrow$
CAGAN (Liu et al. 2021) 9.23 0.685 0.500 0.700 0.722
CAGAN (scaled) 6.83 0.720 0.511 0.894 0.768
Table 4: Ablation study on the CAGAN with FFHQ. “scaled” means that SVS is applied to the pruned weights.
There are only a few pruning methods for generative models, with CAGAN (Liu et al. 2021), DCP-GAN, and Diff-Prune considered key baselines. We extend our analysis to CAGAN to further validate the applicability of SVS beyond state-of-the-art methods. The pruned weights of CAGAN exhibit a similar singular value distribution to DCP-GAN, with existence of dominant singular vectors. As shown in Table 4, applying SVS to CAGAN consistently leads to substantial improvements in all metrics, confirming that applicability of our method.
Ablation Study on Different Scaling Functions
Dataset Scaling functions FID $\downarrow$ P $\uparrow$ R $\uparrow$ D $\uparrow$ C $\uparrow$
FFHQ $\log(x+1)$ 6.18 0.732 0.513 0.960 0.789
$|\log(x)|$ 6.32 0.731 0.519 0.975 0.796
$\sqrt{x}$ 6.60 0.719 0.532 0.874 0.778
Church $\log(x+1)$ 5.30 0.723 0.458 0.933 0.809
$|\log(x)|$ 5.43 0.720 0.453 0.916 0.799
$\sqrt{x}$ 5.18 0.719 0.464 0.934 0.805
Table 5: Ablation on the DCP-GAN (StyleGAN2 (small)) with different scaling functions.
While we primarily use the square root function, we also test alternative scaling functions, such as $\log(x+1)$ and $|\log(x)|$, for reducing singular value gaps. Table 5 shows similar performance across all functions, highlighting the importance of addressing singular value gaps. The square root function was chosen for its simplicity, with results suggesting that broader exploration of scaling strategies could further enhance compression performance.
Ablation Study on Different Scaling Methods
Method (StyleGAN2 (small), FFHQ) FID $\downarrow$
(a) Normalizing singular values N/A
(b) Spectral normalizing singular values N/A
(c) Squaring singular values N/A
(d) Spectral normalization to the generator 8.00
(e) Iterative refining the generator’s weights 8.40
Singular Value Scaling (SVS) 6.60
Table 6: Ablation on the different refinement methods. “N/A” indicates that the model diverges.
In Table 6, we explore additional approaches to refine singular values: (a) normalizing singular values (setting $\sigma_{i}=1$), (b) normalizing with the spectral norm (dividing all singular values $\sigma_{i}$ by the largest singular value $\sigma_{1}$), (c) Squaring singular values, (d) applying spectral normalization to the synthesis network (Miyato et al. 2018), and (e) iterative refining the generator’s weights with our method. For (e), we perform iterative refining only for the first half of the training due to performance degradation. From (a) and (b), we find that setting singular values to 1 (where the weights are orthogonal) or spectral normalizing singular values cause the training to diverge, similar to previous observations (Saxe, McClelland, and Ganguli 2014; Wang and Fu 2023). From (c), we find that widening singular value gaps leads to early training divergence, highlighting the importance of minimizing these gaps for effective fine-tuning. From (d), we see that applying spectral normalization to the generator’s weights makes it difficult for the generator to learn. Finally, from (e), we find that repeatedly refining the singular values of weights during training perturbs the learning process, depriving the generator of sufficient training time and leading to suboptimal results. From these results, we find that applying our method once at model initialization is not only efficient but also leads to more effective training.
Conclusion
In this paper, we propose “Singular Value Scaling” that refines pruned weights to achieve more effective and efficient generative model compression. Our method minimizes the disparities between singular values of pruned weights, allowing the pruned model to explore a broader weight space. Our method is not only simple and effective but also versatile, applicable across different model types, GANs and Diffusion models. Extensive experiments demonstrate that our method provides a more effective initialization state, leading to performance improvements across various metrics for both StyleGAN and Diffusion model compression.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.2020-0-01336, Artificial Intelligence Graduate School Program(UNIST)), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT). (No.2022R1C1C100849612), Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.2022-0-00959, (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making), Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2022-II220264, Comprehensive Video Understanding and Generation with Knowledge-based Deep Logic Neural Network) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-02068, Artificial Intelligence Innovation Hub).